The function f(x)=x+19 is one-to-one.
Find an equation for f^-1(x), the inverse function.
f^-1(x) = ?
Thank you!!!

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

The function f(x)=x+19 is one-to-one.
Find an equation for f^-1(x), the inverse function.
----------
y = x+19
Swap x & y
x = y+19
Solve for y
y = x-19
f^-1(x) = x-19

Answer by greenestamps(13200)



For many relatively simple functions like this one, you can find the inverse by knowing that an inverse function "gets you back where you started".

The given function adds 19 to the input value.

The inverse function gets you back where you started by doing the opposite operations in the opposite order.

In this example, the given function has only one operation, addition. The opposite of addition is subtraction.

So the inverse function subtracts 19 from the input value.

Answer: the inverse function is f^-1(x) = x-19.